OpenStudy (anonymous):

solve by factoring (i suck at this) 3y^2+y-14=0

OpenStudy (anonymous):

im sick and cant think..help please :)

OpenStudy (anonymous):

(y-2)(3y+7)=0

OpenStudy (anonymous):

wat are the solutions?

OpenStudy (anonymous):

y-2=0 3y+7=0

OpenStudy (anonymous):

...woodnt (y-2)= y=2?

OpenStudy (anonymous):

omg its true

OpenStudy (anonymous):

huh? wats true?

OpenStudy (anonymous):

y-2=0 2-2=0 so y=2 isnt' it?

OpenStudy (anonymous):

oh..ya, thts wat i said earlier, but wadda bout the other solution? wood tht be y=-7 or y=-3y?

OpenStudy (anonymous):

3y+7=0 lets substract -7, it's good idea isn't it?

OpenStudy (anonymous):

okay... so its 3y=-7, then u divide rite? so y=-2.333333 rite?

OpenStudy (anonymous):

yeah but you better write -7/3 which is \(-2\frac{1}{3}\) at least i was taught not to approximate, i don't know how you do it in your country, probably US?

OpenStudy (anonymous):

yep..in the U.S, but it wood make more sense to not approximate it, thnx for the help :)